For a given set of data m3 = 10.8 and m2 = 9.2; determine and explain the skewness of the distribution.
25. A given distribution has m4 = 108.8 and m2 = 8.0; determine and explain kurtosis of the distribution
1
Expert's answer
2020-07-22T18:58:55-0400

"24) coefficient\; of \;skewness\; =\frac{m_{3}^2}{m_{2}^3}=\frac{10.8^2}{9.2^3}\approx 0.15\\
\text{the distribution is approximately symmetric.}\\
 25)coefficient\; of \;kurtosis \; =\frac{m_{4}}{m_{2}^2}=\frac{108.8}{8^2}\approx 1.7\\
\text{the distribution is platykurtic.}\\"